OpenStudy (kirbykirby):

If f(x) = integral [cos(t^2)dt] from 0 to x, using the known power series for cos x, find a power series for f(1).

OpenStudy (kirbykirby):

This is f(x): \[f(x)=\int\limits_{0}^{X}\cos(t^2)dt\]

OpenStudy (kirbykirby):

I ended up finding \[\sum_{n=0}^{infinity}\frac{(-1)^{n}x^{4n+1}}{(4n+1)(2n)!}=x-\frac{x^5}{5*2!}+\frac{x^9}{9*4!}-\frac{x^{13}}{13*6!}+...\]

OpenStudy (kirbykirby):

But I'm not quite sure what it means to find the power series for f(1).. Isn't that just a number (whatever I found evaluated at 1)? There is no more x-dependence so how is it a power series :S I'm so confused

OpenStudy (anonymous):

Did you mean using known McLaurin series, or Taylor series for Cos (x)?

OpenStudy (kirbykirby):

No. This is the exact question, and the power series to find is for that function above (the integral one), although the question states to find it for f(1) specifically.